# Kalb-Ramond field, black holes and black strings in (2 + 1)D

### v3

@@
New rotating dilaton black hole and black string solutions in three spacetime dimensions are obtained. The background spacetime interpolates between Anti-de Sitter and a (an asymptotically) flat spacetime. The new black strings are characterized by their masses, angular momenta and axion charges. The uncharged black string solutions have only a single horizon. They contain a curvature singularity enclosed inside their horizons. Thus, they are also black hole solutions of Einstein-scalar gravity. On the other hand, the charged black string solutions have two horizons. They may or may not contain a curvature singularity depending on the ratio of their inner and outer horizons radii. When they contain a singularity, the singularity is either at or enclosed inside their inner horizons. We also obtain, in addition to the uncharged black strings, other new black hole solutions of Einstein-scalar gravity by applying duality transformation on the new charged black strings. The new black holes are described by their masses and angular momenta. We show that their angular momenta do not depend on their horizons radii. We also find that the masses of a class of the black holes can be made negative by adjusting, using the gauge ambiguity, the asymptotic value of the (independent component of the) Kalb-Ramond field in the dual black strings. We also discuss black hole and black string solutions with a curvature singularity at or beyond their outer or event horizons.
@@ 2. BTZ black hole
The metric is invariant under $\rho\to-\rho$. The BTZ black hole metric is obtained after identifying $\omega$ periodically. In the metric (2.6) we now identify $\omega$ with period $2\pi$, i.e., $\omega\sim\omega+2\pi$. This in turn implies $(\overline{\varphi},\overline{\psi})\sim(\overline{\varphi}+2\pi\rho_{+}/l,\overline{\psi}+2\pi\rho_{-}/l)$. Now $\rho$ is radial coordinate, i.e. $\rho\geq 0$. Therefore, BTZ black hole metric can be viewed as a quotient space of $AdS_{3}$. Thus, locally it is isometric to $AdS_{3}$. Furthermore, the scalar curvature is constant. It is given by8 In $AdS_{d}$ the scalar curvature and cosmological constant are given by $R=-{d(d-1)\over l^{2}},\quad\Lambda=-{(d-1)(d-2)\over 2l^{2}}.$ (2.7)2.7( 2.7 )
@@ 3.1. New black string solutions
Other values of $\gamma$ can be mapped onto the interval (3.6) by rescaling the coordinates and the string coupling. The transformation rules under $\gamma\to 1/\gamma$ are $\varphi\to\varphi/\gamma$, $\psi\to-\psi/\gamma$ and $g_{s}^{2}\to g_{s}^{2}/\gamma^{2}$ (see Appendix E in [1]). The metric, the three form antisymmetric tensor and dilaton are invariant by these transformations.
$AdS_{3}$ (with Kalb-Ramond two form) is an exact solution to (classical) string theory, i.e., there are no $\alpha^{\prime}=l_{s}^{2}$ corrections to $AdS_{3}$. $l_{s}$ is the string length. The solution ${\cal A}_{3}$ is related to $AdS_{3}$ by T-dualities and coordinate shift (see [1] for details), and thus it is exact in $\alpha^{\prime}$. However, it is perturbative in $g_{s}$ [19]. On the world-sheet the deformation corresponds to an exact marginal deformation (see, e.g., [1]). More on this in a future work. In the low energy (super)gravity theory the metric, three form flux and dilaton are governed by the action
@@ 3.2. Negative coupling
The global structure of the solutions which satisfy the inequality (3.20) is described using Penrose diagram in Fig. 1. The solutions are axially symmetric. Therefore, we set the (shifted) angular coordinate (in each patch) to a particular value, see Appendix D for details. Thus, a point in the diagram represents a circle in spacetime. The diagram extends infinitely in both vertical directions.
@@
We note that at $\gamma=-1$ we have $\rho_{0}=\rho_{-}$. Thus, at $\gamma=-1$ the inner horizon becomes singular and the black string solution has only one event horizon. Since $\rho_{-}<\rho_{+}$, the curvature singularity is enclosed inside the event horizon. In general, for negative values of $\gamma$, i.e., $-1\leq\gamma\leq(\rho_{-}-\rho_{+})/(\rho_{-}+\rho_{+})$, we have $0\leq\rho_{0}\leq\rho_{-}$. $\rho_{0}=0$ when $\gamma=(\rho_{-}-\rho_{+})/(\rho_{-}+\rho_{+})$. Thus, the singularity is at or enclosed inside the inner horizon.
@@
The Ricci scalar is singular at the origin.16 Thus, here $x$ need not be periodic. Although we assumed $\rho_{+}>\rho_{-}$, i.e. $\gamma\neq 0$, to arrive at the solution, also here the limit $\rho_{+}\to\rho_{-}$ is mathematically well defined and gives non-trivial solution. We do not consider the limit here. We also note that both the dilaton and Kalb-Ramond fields are singular at the origin. We will discuss this family of solutions further later in the section. We in particular show that they are related to the solutions discussed in [21]. The global structure of the solutions for which $0\leq\rho_{0}<\rho_{-}$ is described using Penrose diagram in Fig. 2. The diagram is obtained in the same way as the example in Appendix D. In the diagram we set, in each patch, the (shifted) angular coordinate to a particular value since any other value is equivalent and related by axial symmetry. Therefore, a point in the diagram represents a circle in spacetime.
@@
The curvature singularity is at the origin. This solution in the gauge in which $B=0$ is dual to the non-rotating BTZ black hole with a non-vanishing $B$ [20], see also [22].17 As we will show later in the paper, for a general gauge, it is also equivalent to a non rotating black hole in a spacetime with non-vanishing angular velocity at radial infinity and no $B$ field. It is the product of the two dimensional black hole [[27]27,,  [28]28] with $S^{1}$. To see this define
@@ 3.3. Positive coupling
Thus, for an observer residing in the spacetime region $\rho<\rho_{-}$, i.e., the exterior region, the solution (3.33) has a singularity at $\rho=\rho_{0}=\rho_{+}$ beyond the horizon, i.e., $\rho_{0}>\rho_{-}$. Thus, in view of this observer, (3.33) describes a black hole. Since $x$ is compact, the solution contains closed time-likes curve in the region $\rho>\rho_{0}=\rho_{+}$. Thus, they are located in the interior region. We note that the $B$ field, i.e. $B_{tx}$, is a non-zero constant. We can fix the $B$ field to zero using the gauge freedom. The Penrose diagram is given in Fig. 5.
@@ 3.4. New black hole solutions
$\eqalign{d{\overline{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over\left(1-{r_{+}\over r}\right)\left(1-{r_{-}\over r}\right)}-\left(1-{r_{+}\over r}\right)d{\tau}^{2}+\left(1-{r_{-}\over r}\right)d{\chi}^{2}\right],\cr B&=B_{{\tau}{\chi}}d{\tau}\wedge d{\chi},\quad B_{{\tau}{\chi}}={l^{2}\over 4}\left({\sqrt{r_{-}r_{+}}\over r}-1\right),\cr-2\Phi&=\ln\left({|r|\over r_{+}}\right)+{\tenrm const.},}$ (3.42)3.42( 3.42 )
where $r_{\pm}=\rho_{\pm}^{2}$.18 The constant is ${\tenrm const.}=\ln\left[(1-\gamma)^{2}/g_{s}^{2}\right]$. It is invariant under $\gamma\to 1/\gamma$. Note $\Phi_{\tenrm there}=-2\Phi_{\tenrm here}$. Note also in general $\chi$ need not be periodic and $r$ can take negative values. However here we will assume, for simplicity, $\chi$ is compact and $r$ is positive. The Ricci scalar is singular at $r=0$ and it is finite at $r=r_{\pm}$. We note that inside the event horizon, i.e., in the region $0<r<r_{+}$, the coordinate $\tau$ is everywhere space-like. We also note that $\chi$ becomes time-like in the region $0<r<r_{-}$. We will not discuss (3.42) further here. We refer to [21] for more details and a comprehensive discussion.
@@
where ${\widetilde{x}}$ is the dual coordinate. Under $\gamma\to 1/\gamma$, ${\widetilde{x}}\to-{\widetilde{x}}/\gamma$ and ${\widehat{t}}\to{\widehat{t}}/\gamma$. We note that the $\widetilde{B}$ field, and thus in turn the gauge parameter $\lambda$, enters into the coefficient of the cross term $d{\widehat{t}}d{\widetilde{x}}$. We will assume $0<\lambda<-1/\gamma$ so that the solution describes a black hole. The solution is invariant under $\gamma\to 1/\gamma$ because $\lambda$ transforms as $\lambda\to-\gamma(1+\gamma\lambda)$. The scalar curvature is given by
@@
Thus, the black hole solutions have a curvature singularity enclosed inside their event horizons.
@@ 4. Dilatonic rotating black hole solutions
the quasi-local energy $E$, angular momentum $j$ and mass $m$ are given by (with $\kappa^{2}=\pi$)
@@
The (analogous) ADM mass $M$ and angular momentum $J$ are then given by
@@
This solution (4.26) gives $\varepsilon=\sqrt{-\Lambda\rho^{2}}$. The particular solution is chosen such that at constant time or $\varphi$ it asymptotically coincides or merges at large $\rho$ smoothly with the solution (2.6) [[3]3,,[4]4]. Thus, the quasi-local mass is given by
@@
$E={\sqrt{\rho_{+}^{2}-\rho_{-}^{2}}\over l}\left[2-\left({2\rho^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{-}^{2}}\right)\sqrt{1-{\rho_{+}^{2}\over\rho^{2}}}\right],\quad\rho\geq\rho_{+}.$ (4.35)4.35( 4.35 )
@@
Similarly, for the black hole solutions (4.16) we find that the angular momenta are given by20 Here we set $g_{s}=1$.
@@
Although non-zero $\Omega$ can certainly lower (and raise) the mass, see e.g. the general expression (C.27) of the mass in the appendix, it does not always result in negative mass solutions. A simple example is the BTZ solution with non vanishing rotation at asymptotic infinity. To allow rotation at infinity we shift $Z$ in (4.29) by a non-zero constant. We denote the constant by $\overline{\lambda}$ and make the change
@@ 5. Discussion
The coupling $\gamma$ takes its value in the range $[-1,1]$. The spacetime ${\cal A}_{3}$ (5.1) interpolates between $AdS_{3}$ and (asymptotically) flat spacetime ${\tenrm I\kern-1.79993ptR}\times S^{1}\times{\tenrm I\kern-1.79993ptR}$. In [1] the deformation is interpreted in general as moving holographic boundaries. The spacetime ${\cal A}_{3}$ is free of curvature singularity when $\gamma$ is negative. Thus, our discussion has been restricted mainly to the case $-1\leq\gamma\leq 0$. For positive $\gamma$ the deformation generates a curvature singularity from the onset [1].
The black string solutions are described by (3.15) which we again write below for convenience. They carry the axion charge per unit length $Q$ (3.19). The black string solutions are
@@
$\lambda$ is an arbitrary constant which determines the value of the $B$ field, i.e. $B_{tx}$, at $\rho=\infty$. It does not enter in $H$. However, it has a non trivial implication. The black string solution (5.3) has inner and outer horizons. This is also a general feature of rotating black holes. The inner horizon is at $\rho=\rho_{-}$ and the outer horizon is at $\rho=\rho_{+}$. In the case $x$ is compact and there is no curvature singularity we identify it so that there is no conical singularity. Its period does not depend on the gauge parameter $\lambda$. In this paper, however, we in general assume $x$ is compact and $\rho$ is positive. To the best of my knowledge the black string solutions (5.3) have not been reported elsewhere in the literature.
@@
It follows from the cases $\gamma=\pm 1$ that both inner and outer horizons can be unstable upon coupling rotating black holes to matter.
We Lorentz boost the black string solution (5.3) along the $x$ direction with a specific rapidity (3.36),
@@
The Kalb-Ramond field is invariant under Lorentz boost. The metric however now has no cross term and thus static, see (3.37). Applying T-duality transformation along the $\widehat{x}$ direction gives the black hole solution (3.44),
@@
We note that the black hole depends on the gauge parameter $\lambda$. In particular we can have black holes with negative masses (4.44). The mass is positive if we choose $-1/2\gamma\leq\lambda<-1/\gamma$. Note in the case of BTZ solutions, those solutions with negative masses do not correspond to black holes, see (2.13) and (2.14). We also note that the deformation slows down the rotation [1]. The norm of $J$ decreases as we increase the norm of $\gamma$. The axion charge per unit length $Q$ of the black string (with $g_{s}=1$) equals the angular momentum $J$ of the black hole [22]. At $\lambda=-1/2\gamma$, the angular velocity at infinity $\Omega=0$. In this case the ADM mass and angular momentum reduce to (4.43)
@@
At $\gamma=-1$ we note that $J=0$ and the mass $M$ (5.10) is independent of the gauge parameter $\lambda$. $M$ is determined by the radii $\rho_{+}$ and $\rho_{-}$.
In this paper we also noted that a solution with a non-zero asymptotic Kalb-Ramond field is equivalent to a solution with no Kalb-Ramond field and no vanishing rotation at asymptotic infinity.
@@
We also hope to study other string solutions by gluing those solutions of positive $\gamma$ with negative $\gamma$. We give below an example which is obtained by gluing (3.25) and (3.33). It is convenient to introduce first the parameters $0<{\overline{\rho}}_{-}<{\overline{\rho}}_{+}=\rho_{-}<\rho_{+}$. We define $\alpha=(\rho_{+}-{\overline{\rho}}_{-})/\rho_{-}$. We assume $\rho_{+}<\sqrt{2}\rho_{-}$. We furthermore assume ${\overline{\rho}}_{-}=\sqrt{2\rho_{-}^{2}-\rho_{+}^{2}}$.
@@
$\eqalign{ds^{2}=&{l^{2}\rho^{2}d\rho^{2}\over({\overline{\rho}}_{+}^{2}-\rho^{2})({\overline{\rho}}_{-}^{2}-\rho^{2})}-{({\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2})({\overline{\rho}}_{-}^{2}-\rho^{2})dt^{2}\over 4l^{2}\rho^{2}}+{({\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2})\rho^{2}\over 4({\overline{\rho}}_{+}^{2}-\rho^{2})}\left[dx-{1\over l}\left(\alpha+{{\overline{\rho}_{+}}{\overline{\rho}_{-}}\over\rho^{2}}\right)dt\right]^{2},\cr B_{tx}&={{\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2}\over 4l}={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&={g_{s}^{2}\over 4}\left({\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2}\over{\overline{\rho}}_{+}^{2}-\rho^{2}\right).}$ (5.13)5.13( 5.13 )
@@
The solution has the peculiar feature that it contains a ring curvature singularity at $\rho={\overline{\rho}}_{+}=\rho_{-}$ in between the inner and outer horizons. The inner horizon is at $\rho={\overline{\rho}}_{-}$. The outer horizon is at $\rho=\rho_{+}$. Note they are not related by sign reflection, i.e. ${\overline{\rho}}_{-}\neq-\rho_{+}$. The Penrose diagram of the solution is given in Fig. 6. It is obtained by combining Fig. 3 and 5.
@@
We hope to study the corrections in $g_{s}$ and their roles in resolving the singularity in a future work.

### v4

@@
New rotating dilaton black hole and black string solutions in three spacetime dimensions are obtained. The background spacetime interpolates between Anti-de Sitter and a (an asymptotically) flat spacetime. The new black strings are characterized by their masses, angular momenta and axion charges. The uncharged black string solutions have only a single horizon. They contain a curvature singularity enclosed inside their horizons. Thus, they are also black hole solutions of Einstein-scalar gravity. On the other hand, the charged black string solutions have two horizons. They may or may not contain a curvature singularity depending on the ratio of their inner and outer horizons radii. When they contain a singularity, the singularity is either at or enclosed inside their inner horizons. We also obtain, in addition to the uncharged black strings, other new black hole solutions of Einstein-scalar gravity by applying duality transformation on the new charged black strings. The new black holes are described by their masses and angular momenta. We show that their angular momenta do not depend on their horizons radii. We also find that the masses of a class of the black holes can be made negative by adjusting, using the gauge ambiguity, the asymptotic value of the (independent component of the) Kalb-Ramond field in the dual black strings. We also discuss black hole and black string solutions with a curvature singularity at or beyond their outer or event horizons. Black hole solutions with a ring curvature singularity in between their inner and outer horizons are also presented.
@@ 2. BTZ black hole
The metric is invariant under $\rho\to-\rho$. The BTZ black hole metric is obtained after identifying $\omega$ periodically. In the metric (2.6) we now identify $\omega$ with period $2\pi$, i.e., $\omega\sim\omega+2\pi$. This in turn implies $(\overline{\varphi},\overline{\psi})\sim(\overline{\varphi}+2\pi\rho_{+}/l,\overline{\psi}+2\pi\rho_{-}/l)$. Now $\rho$ is radial coordinate, i.e. $\rho\geq 0$. Therefore, BTZ black hole can be viewed as a quotient space of $AdS_{3}$. Thus, locally it is isometric to $AdS_{3}$. Furthermore, the scalar curvature is constant. It is given by8 In $AdS_{d}$ the scalar curvature and cosmological constant are given by $R=-{d(d-1)\over l^{2}},\quad\Lambda=-{(d-1)(d-2)\over 2l^{2}}.$ (2.7)2.7( 2.7 )
@@ 3.1. New black string solutions
Other values of $\gamma$ can be mapped onto the interval (3.6) by rescaling the coordinates and the string coupling at $\gamma=0$, i.e., $g_{s}$. The transformation rules under $\gamma\to 1/\gamma$ are $\varphi\to\varphi/\gamma$, $\psi\to-\psi/\gamma$ and $g_{s}^{2}\to g_{s}^{2}/\gamma^{2}$ (see Appendix E in [1]). The metric, the three form antisymmetric tensor and dilaton are invariant by these transformations.
$AdS_{3}$ (with Kalb-Ramond two form) is an exact solution to (classical) string theory, i.e., there are no $\alpha^{\prime}=l_{s}^{2}$ corrections to $AdS_{3}$. $l_{s}$ is the string length. The solution ${\cal A}_{3}$ is related to $AdS_{3}$ by T-dualities and coordinate shift (see [1] for details), and thus it is exact in $\alpha^{\prime}$. However, it is perturbative in $g_{s}$ [19]. On the world-sheet the deformation corresponds to an exact marginal deformation (see, e.g., [1]). More on this in a future work. Later in the section we also give previously known $\alpha^{\prime}$ exact (black string) solutions (in ${\cal A}_{3}$) as special cases.
In the low energy (super)gravity theory the metric, three form flux and dilaton are governed by the action
@@ 3.2. Negative coupling
The global structure of the solutions which satisfy the inequality (3.20) is described using Penrose diagram in Fig. 1. The solutions are axially symmetric. Therefore, we set the (shifted) angular coordinate (in each patch) to a particular value, see Appendix D for details. Thus, a point in the diagram represents a circle in spacetime, i.e. the orbit of the axial isometry. The diagram extends infinitely in both vertical directions.
@@
We note that at $\gamma=-1$ we have $\rho_{0}=\rho_{-}$. Thus, at $\gamma=-1$ the inner horizon becomes singular and the black string solution has only one event horizon. Since $\rho_{-}<\rho_{+}$, the curvature singularity is enclosed inside the event horizon. In general, for negative values of $\gamma$ and $\rho_{-}/\rho_{+}\geq\mu$, i.e., $-1\leq\gamma\leq(\rho_{-}-\rho_{+})/(\rho_{-}+\rho_{+})$, we have $0\leq\rho_{0}\leq\rho_{-}$. $\rho_{0}=0$ when $\gamma=(\rho_{-}-\rho_{+})/(\rho_{-}+\rho_{+})$. Thus, the singularity is at or enclosed inside the inner horizon.
@@
The Ricci scalar is singular at the origin.16 Thus, here $x$ need not be periodic (with a definite period to avoid a conical singularity). Although we assumed $\rho_{+}>\rho_{-}$, i.e. $\gamma\neq 0$, to arrive at the solution, also here the limit $\rho_{+}\to\rho_{-}$ is mathematically well defined and gives non-trivial solution. In this limit, after an appropriate change of coordinates, we obtain $ds^{2}=(l^{2}/4)\left[dr^{2}/(r-r_{0})^{2}-d\tau^{2}(1-(r_{0}/r))^{2}+(d\chi-(r_{0}/r)d\tau)^{2}\right]$, $B=B_{\tau\chi}d\tau\wedge d\chi,B_{\tau\chi}=-(l^{2}/4)(1-(r_{0}/r))$, $e^{2\Phi}=g_{s}^{2}r_{0}/r$, where $r_{0}=\rho_{+}^{2}=\rho_{-}^{2}$. The scalar curvature is $R=\left[2(4r-7r_{0})r_{0}\right]/(l^{2}r^{2})$. Thus, $r=r_{0}$ is not a curvature singularity. We do not discuss this solution further here. We also note that both the dilaton and Kalb-Ramond fields are singular at the origin. We will discuss this family of solutions further later in the section. We in particular show that they are related to the solutions discussed in [21]. The global structure of the solutions for which $0\leq\rho_{0}<\rho_{-}$ is described using Penrose diagram in Fig. 2. The diagram is obtained in the same way as the example in Appendix D. In the diagram we set, in each patch, the (shifted) angular coordinate to a particular value since any other value is equivalent and related by axial symmetry. Therefore, a point in the diagram represents a circle in spacetime.
@@
The curvature singularity is at the origin. This solution in the gauge in which $B=0$ is dual to the non-rotating BTZ black hole with a non-vanishing $B$ [20], see also [22].17 As we will show later in the paper, for a general gauge, it is also equivalent to a non-rotating black hole in a spacetime with non-vanishing angular velocity at radial infinity and no $B$ field. It is the product of the two dimensional black hole [[27]27,,  [28]28] with $S^{1}$. To see this define
@@ 3.3. Positive coupling
Thus, for an observer residing in the spacetime region $\rho<\rho_{-}$, i.e., the exterior region, the solution (3.33) has a singularity at $\rho=\rho_{0}=\rho_{+}$ beyond the horizon, i.e., $\rho_{0}>\rho_{-}$. Thus, in view of this observer, (3.33) describes a black hole. Since $x$ is compact, the solution contains closed time-like curves in the region $\rho>\rho_{0}=\rho_{+}$. Thus, they are located in the interior region. We note that the $B$ field, i.e. $B_{tx}$, is a non-zero constant. We can fix the $B$ field to zero using the gauge freedom. The Penrose diagram is given in Fig. 5.
@@ 3.4. New black hole solutions
$\eqalign{d{\overline{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over r^{2}\left(1-{r_{+}\over r}\right)\left(1-{r_{-}\over r}\right)}-\left(1-{r_{+}\over r}\right)d{\tau}^{2}+\left(1-{r_{-}\over r}\right)d{\chi}^{2}\right],\cr B&=B_{{\tau}{\chi}}d{\tau}\wedge d{\chi},\quad B_{{\tau}{\chi}}={l^{2}\over 4}\left({\sqrt{r_{-}r_{+}}\over r}-1\right),\cr-2\Phi&=\ln\left({|r|\over r_{+}}\right)+{\tenrm const.},}$ (3.42)3.42( 3.42 )
where $r_{\pm}=\rho_{\pm}^{2}$.18 The constant is ${\tenrm const.}=\ln\left[(1-\gamma)^{2}/g_{s}^{2}\right]$. It is invariant under $\gamma\to 1/\gamma$. Note $\Phi_{\tenrm there}=-2\Phi_{\tenrm here}$. Note also in general $\chi$ need not be periodic (with a definite period) and $r$ can take negative values. However here we will assume, for simplicity, $\chi$ is compact and $r$ is positive. The Ricci scalar is singular at $r=0$ and it is finite at $r=r_{\pm}$. We note that inside the event horizon, i.e., in the region $0<r<r_{+}$, the coordinate $\tau$ is everywhere space-like. We also note that $\chi$ becomes time-like in the region $0<r<r_{-}$. We will not discuss (3.42) further here. We refer to [21] for more details and a comprehensive discussion.
@@
where ${\widetilde{x}}$ is the dual coordinate. Under $\gamma\to 1/\gamma$, ${\widetilde{x}}\to-{\widetilde{x}}/\gamma$ and ${\widehat{t}}\to{\widehat{t}}/\gamma$. We note that the $\widetilde{B}$ field, and thus in turn the gauge parameter $\lambda$, enters into the coefficient of the cross term $d{\widehat{t}}d{\widetilde{x}}$. The solution is invariant under $\gamma\to 1/\gamma$ because $\lambda$ transforms as $\lambda\to-\gamma(1+\gamma\lambda)$. We will assume $0<\lambda<-1/\gamma$ so that the solution describes a black hole. The scalar curvature is given by
@@
Thus, the black hole solutions (labelled by the triplet $(\rho_{-},\rho_{+};\lambda)$) have a curvature singularity enclosed inside their event horizons.
@@ 4. Dilatonic rotating black hole solutions
the quasi-local energy $E$, angular momentum $j$ and mass $m$ are given by (with $\kappa^{2}=\pi$ in some length unit)
@@
The angular momentum $J$ and (analogous) ADM mass $M$ are then given by
@@
This solution (4.26) gives $\varepsilon=\sqrt{-\Lambda\rho^{2}}$. The particular solution is chosen such that at constant time or $\varphi$ it asymptotically coincides or merges at large $\rho$ smoothly with the solution (2.6) [[3]3,,[4]4]. Thus, the quasi-local mass $m$ is given by20 Note here the boundary $B_{t}$ is at large $\rho$, see Appendix C. Thus, in the expressions for $m$ and $E$, $\rho$ is assumed to be large. The same applies for the quasi-local masses and energies that appear below. Here they serve as intermediate results.
@@
$E={\sqrt{\rho_{+}^{2}-\rho_{-}^{2}}\over l}\left[2-\left({2\rho^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{-}^{2}}\right)\sqrt{1-{\rho_{+}^{2}\over\rho^{2}}}\right],\quad\rho>\rho_{+}.$ (4.35)4.35( 4.35 )
@@
Similarly, for the black hole solutions (4.16) we find that the angular momenta are given by21 Here we set $g_{s}=1$.
@@
Although non-zero $\Omega$ can certainly lower (and raise) the mass, see e.g. the general expression (C.27) of the mass in the appendix, it does not always result in negative mass solutions. A simple example is the BTZ solution with non-vanishing rotation at asymptotic infinity. To allow rotation at infinity we shift $Z$ in (4.29) by a non-zero constant. We denote the constant by $\overline{\lambda}$ and make the change
@@ 5. Discussion
The coupling $\gamma$ takes its value in the range $[-1,1]$. The spacetime ${\cal A}_{3}$ (5.1) interpolates between $AdS_{3}$ and (asymptotically) flat spacetime ${\tenrm I\kern-1.79993ptR}\times S^{1}\times{\tenrm I\kern-1.79993ptR}$. It is obtained by applying a sequence of $\alpha^{\prime}$ exact transformations on $AdS_{3}$ [1]. In [1] the deformation is interpreted in general as moving holographic boundaries. The spacetime ${\cal A}_{3}$ is free of curvature singularity when $\gamma$ is negative. Thus, our discussion has been restricted mainly to the case $-1\leq\gamma\leq 0$. For positive $\gamma$ the deformation generates a curvature singularity from the onset [1].
The black string solutions are described by (3.15) which we again write below for convenience. They are obtained from ${\cal A}_{3}$ by applying the simple coordinate transformation (3.13) (and making the change of variable (3.14)). They carry the axion charge per unit length $Q$ (3.19). The black string solutions are
@@
$\lambda$ is an arbitrary constant which determines the value of the $B$ field, i.e. $B_{tx}$, at $\rho=\infty$. It does not enter in $H$. However, it has a non-trivial implication. The black string solution (5.3) has inner and outer horizons. This is also a general feature of rotating black holes. The inner horizon is at $\rho=\rho_{-}$ and the outer horizon is at $\rho=\rho_{+}$. In this paper, we in general assume, unless stated otherwise, $x$ is compact and $\rho$ is positive. In the case there is no curvature singularity we identify $x$ so that there is no conical singularity. Its period does not depend on the gauge parameter $\lambda$. To the best of my knowledge the black string solutions (5.3) have not been reported elsewhere in the literature.
@@
It follows from the cases $\gamma=\pm 1$ that both inner and outer horizons can be unstable upon coupling rotating black holes to matter. See below also for a novel case in which a ring singularity forms in between the inner and outer horizons.
In section three, we Lorentz boosted the black string solution (5.3) along the $x$ direction with a specific rapidity (3.36),
@@
The Kalb-Ramond field is invariant under the Lorentz boost. The resulting metric however now has no cross term and thus static, see (3.37). We subsequently applied T-duality transformation along the $\widehat{x}$ direction. This resulted in the black hole solution (3.44),
@@
We note that the black hole depends on the gauge parameter $\lambda$. In particular we can have black holes with negative masses (4.44). The mass is positive if we choose $-1/2\gamma\leq\lambda<-1/\gamma$. Note in the case of BTZ solutions, those solutions with negative masses do not correspond to black holes, see (2.13) and (2.14). We also note that the deformation slows down the rotation [1]. That is, the norm of $J$ decreases as we increase the norm of $\gamma$. The axion charge (per unit length) $Q$ of the black string (with $g_{s}=1$) equals the angular momentum $J$ of the black hole. This is consistent with the work [22].
At $\gamma=-1$ we note that $J=0$ and the mass $M$ is independent of the gauge parameter $\lambda$. $M$ is determined by the radii $\rho_{+}$ and $\rho_{-}$. At $\lambda=-1/2\gamma$, the angular velocity at infinity is zero, i.e., $\Omega=0$. In this case the ADM mass and angular momentum reduce to (4.43)
@@
In this paper we also noted that a solution with a non-zero asymptotic Kalb-Ramond field is equivalent to a solution with no Kalb-Ramond field and no vanishing rotation at asymptotic infinity. In general, non-zero rotation allowes negative mass solutions.
@@
We also hope to study other string solutions by gluing those solutions of positive $\gamma$ with negative $\gamma$. We give below an example which is obtained by gluing (3.25) and (3.33). It is convenient to introduce first the parameters $0<{\overline{\rho}}_{-}<{\overline{\rho}}_{+}=\rho_{-}<\rho_{+}$. We define $\delta=(\rho_{+}-{\overline{\rho}}_{-})/\rho_{-}$. We assume $\rho_{+}<\sqrt{2}\rho_{-}$. We furthermore assume ${\overline{\rho}}_{-}=\sqrt{2\rho_{-}^{2}-\rho_{+}^{2}}$.
@@
$\eqalign{ds^{2}=&{l^{2}\rho^{2}d\rho^{2}\over({\overline{\rho}}_{+}^{2}-\rho^{2})({\overline{\rho}}_{-}^{2}-\rho^{2})}-{({\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2})({\overline{\rho}}_{-}^{2}-\rho^{2})dt^{2}\over 4l^{2}\rho^{2}}+{({\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2})\rho^{2}\over 4({\overline{\rho}}_{+}^{2}-\rho^{2})}\left[dx-{1\over l}\left(\delta+{{\overline{\rho}_{+}}{\overline{\rho}_{-}}\over\rho^{2}}\right)dt\right]^{2},\cr B_{tx}&={{\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2}\over 4l}={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&={g_{s}^{2}\over 4}\left({\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2}\over{\overline{\rho}}_{+}^{2}-\rho^{2}\right).}$ (5.13)5.13( 5.13 )
@@
The scalar curvature of the solution (which is described by (5.13) and (5.14)) is $R=-4\xi^{2}\cdot(\rho_{+}^{2}/l^{2})\cdot(1/|\rho^{2}-\rho_{-}^{2}|)$, where $\xi^{2}=1-(\rho_{-}/\rho_{+})^{2}$. The Kretschmann scalar is $K=R^{2}$. Thus, the solution has the peculiar feature that it contains a ring curvature singularity at $\rho={\overline{\rho}}_{+}=\rho_{-}$ in between the inner and outer horizons. The inner horizon is at $\rho={\overline{\rho}}_{-}$. The outer horizon is at $\rho=\rho_{+}$. Note they are not related by sign reflection, i.e. ${\overline{\rho}}_{-}\neq-\rho_{+}$. The Kalb-Ramond field is a constant (field). It can be fixed to zero. The Penrose diagram of the solution is given in Fig. 6. It is obtained by combining Fig. 3 and 5.
@@
Note usually the diagram ends at $\rho={\overline{\rho}}_{+}=\rho_{-}$, however here, since there are no closed time-like curves, the diagram extends outward to infinity and inward to zero. Therefore, the maximal extension simply extends infinitely in both the vertical directions.
We expect the black hole solution (described by (5.13) and (5.14)) to have a generalization in (3 + 1) dimensions with similar feature. That is, with a curvature singularity in between the inner and outer horizons. We hope to study the corrections in $g_{s}$ and their roles in resolving the singularity in a future work.